# Statistical Learning for Heterogeneous Treatment Effects: Pretraining, Prognosis, and Prediction

### v1

@@ 3.1 Simulations under Varying Shared Support Sizes
Our experiments are conducted assuming a randomized trial setting. In this first experiment, we suppose $n$ units, for each we are given a binary treatment assignment $W_{i}\sim Binom(e(X_{i}))$, with $e(X_{i})=0.5$, $p$ covariates $X_{i}\sim\mathcal{N}(0,1)^{p}$ and an outcome $Y_{i}=\mu_{i}+\varepsilon_{i}$. We define $\mu_{i}=\beta_{0}+\sum_{j=1}^{n}X_{ij}\beta_{j}+W_{i}\cdot\tau(X_{i})$, with $\tau(X_{i})=\theta_{0}+\sum_{j=1}^{n}X_{ij}\theta_{j}$, where $\beta_{j}\sim U(0.5,1)$, $\theta_{j}\sim U(1,2)$, and $\varepsilon_{i}\sim\mathcal{N}\left(0,\ \frac{Var(\mu)}{SNR}\right)$. We then introduce some sparsity and limit the size of the active support for $m(X_{i})$ and $\tau(X_{i})$ to $p/3$, respectively. We start out with a fully disjoint support where, which we gradually increase until the support is fully shared, i.e. equal to $p/3$. We evaluate performance in terms of expected mean-squared error (MSE) $\frac{1}{n}\sum_{i=1}^{n}[\tau(X_{i})-\hat{\tau}(X_{i})]^{2}$, which we normalize by the median MSE of the baseline $R$-lasso.
In our first experiment, we looked at the effect of pretraining the $R$-lasso in the two extreme cases, where the support is either fully disjoint or fully shared. Figure 3 shows the performance of the pretrained $R$-lasso in comparison to the baseline $R$-lasso in both settings and with different pretraining strengths (upper panel). Panel b shows the evolution of performance as the shared support increases fully disjoint to fully joint support. In the absence of any overlap, pretraining shows no effect. With increasing shared support, the pretrained $R$-lasso achieves lower MSE, especially under strong pretraining (low $\alpha$).
@@
Figure 4: Concordance of Training Set R-Loss and Test Set MSE. The figure shows the median performance (standardized by the median $R$-lasso without pretraining) oat different pretraining strengths of the $R$-lasso in a linear setting with $n=500$, $p=150$, $SNR=2$, an active set of size $|S|=\frac{p}{3}$ and for $5000$ repetitions. The results are stratified by the size of shared support between prognostic and predictive factors: disjoint (left), partially shared (middle) and fully shared support (right panel). The minimum test set MSE and training set R-loss are concordant. In settings with stronger overlap, stronger pretraining (lower $\alpha$) is favorable. The solid line at 1.0 shows the $R$-lasso without pretraining ($\alpha=1$) for comparison. Error bars denote the interquartile range between the 25th and 75th percentiles.
@@ 3.2 Assessing Power to Detect Heterogeneity
The end goal of fitting CATE models is to discover potential subgroups or strata as defined by our predictor variables $X_{i}$ that have average treatment effects that differ meaningfully from the overall population average. Quantifying their ability to achieve this goal provides valuable information. Recently proposed metrics useful for this task take inspiration from predictive analysis tools such as the ROC curve and include Chernozhukov et al. (2018); Radcliffe (2007); Sverdrup et al. (2024); Imai and Li (2023); Yadlowsky et al. (2025). An intuitive approach to operationalize this is to take inspiration from classical machine learning, where define a training set for fitting a CATE model, and evaluate its performance on a held-out test set. We can then rank individuals from highest to lowest predicted CATE. In the presence of heterogeneity, this should yield a decreasing quantile average treatment effect as we move from higher to lower quantiles. One way to operationalize this CATE-based ranking is to use the targeting operator characteristic (TOC) curve (Yadlowsky et al., 2025), defined as
@@
The area under the TOC (AUTOC) is a numeric measure of heterogeneity and allows us to evaluate whether and how well our CATE model detects heterogeneity against a sharp null of no heterogeneity detected. One limitation of this approach is that train-test splits can be costly and suffer from random split sensitivity, especially in settings with small sample sizes. To overcome this limitation, we adopt a sequential cross-fold approach (Wager, 2024).
Using the TOC curve, we can assess the power at which our CATE model detects heterogeneity. We hypothesized that in the presence of shared support, the pretrained $R$-lasso should be more powerful in detecting heterogeneity than without pretraining. For this experiment, we slightly tweak our DGP to a more challenging setting and reduce the support size to $p/10$, half of which is shared between $m(X_{i})$ and $\tau(X_{i})$ and the rest disjoint. We further introduce correlation among our covariates $X\sim\mathcal{N}_{p}(\mathbf{0},\Sigma),\text{where }\Sigma\text{ is a Toeplitz matrix with entries }\Sigma_{ij}=0.5^{|i-j|}$ and reduce the SNR to 0.5. Figure 5 shows the TOC curve in two configurations: fully disjoint support and partially shared support. In the presence of overlap, the pretrained $R$-lasso shows a superior AUTOC as well as high power in detecting heterogeneity as we vary the the sample size $n$ and the number of covariates $p$ (Figure 6).
@@ 3.3 Sparsity and Reliable Support Recovery: uniR-Lasso
As demonstrated in the previous section, the effectiveness of our pretraining strategy depends both on the presence of overlap and the $m(X_{i})$ model’s ability to accurately detect it. The lasso has shown to be less reliable for this task in the presence of highly correlated futures and false positives in the active set (Su et al., 2017). To address this concern, Chatterjee et al. (2025) recently developed a univariate-guided sparse regression termed “UniLasso” that achieves higher sparsity by constraining the coefficients to have the same sign and reflecting their absolute magnitudes in univariate regression.
Sparsity is also a desirable property in our setting as it will result in more covariates being penalized. Specifically, it will result in stronger, more precise pretraining when fitting the model for $\tau(X_{i})$. In addition, this model also achieves higher interpretability in high-dimensional settings with low signal, which might be desirable in many settings. As highlighted by Candes et al. (2018), identifying a subset of relevant covariates, while controlling false discovery rates (FDR), is highly valuable in many areas (Storey and Tibshirani, 2003). For example in biology, the reliable identification of a single driver of heterogeneity can help discover a larger regulatory network that modifies treatment effects.
@@ 4 Leveraging Side Information in Non-Linear Settings
In many applications, the mean outcome and CATE functions might involve complex interactions and non-linearities that the linear lasso outlined in section 3 may not capture. A straightforward solution to continue using the lasso without modifying Algorithm 1 is to substitute the feature matrix $X_{i}$ with $\psi(X_{i})$ where $\psi:\mathcal{X}\mapsto\mathbb{R}^{d}$ is some some featurization that maps the covariates to a richer $d$-dimensional space. $\psi$ could, for example, be a collection of polynomial basis function. With a rich set of basis functions this approach can sufficiently capture complicated CATE or mean outcome functions (e.g., Chen (2007)). While a wide variety of basis expansions exist and the optimal choice is often application-specific (Friedman, 1991), here we concentrate on gradient boosted trees to generate these basis expansions (Chen and Guestrin, 2016; Friedman, 2001) and use the statistical software package xgboost (Chen and Guestrin, 2016). Gradient boosting is an additive model that fits a sequence of weighted trees or weak learners. We explore its use in the $R$-learner for $M=100$ and $500$ boosting rounds and a tree depth of 1 (stump). This results in a total of $M$ trees (and thus new features), each of which performs a single split. As we increase the tree depth and number of boosting rounds, the transformed feature space is able to capture more complex relations, but also suffers from higher dimensions. To perform hyperparameter tuning, i.e. select the strength of pretraining and the number of boosting rounds, we use the $R$-loss. By concatenating this new feature matrix with our original one, the lasso can now pick from the original and transformed features. This is particularly useful in partially linear settings, where estimating $m(X_{i})$ and $\tau(X_{i})$ might require a model capable of handling both linear and non-linear components simultaneously. Depending on whether we expect $m(X_{i})$ and $\tau(X_{i})$ to differ substantially in their functional forms, we can generate distinct sets of basis functions tailored to each prediction task in the $R$-learner.
@@
The lasso solution to sharing side-information leverages the lasso’s adaptive regularization feature, but similar logic can be applied to other learners, such as random forests. Causal forests (Athey et al., 2019) are forest-based instantiations of the $R$-learner that uses regression forests for $m(X_{i})$ and $e(X_{i})$, and then fits $\tau(X_{i})$ using a forest-weighted version of (3) designed to express heterogeneity in the CATE. As trees rely on split criteria rather than shrinking coefficients as in the lasso, we require a new strategy that allows us to leverage “side information”. An effective way to improve random forests is to use adaptive feature-weighting (e.g., Basu et al. (2018)). Adapted to our setting, this procedure consists of fitting a random forest for $m(X_{i})$, deriving the feature importance score for each covariate, and, finally, sampling each feature in $\tau(X_{i})$ with a probability that reflects its importance in $m(X_{i})$. As for the lasso, we can introduce a hyperparameter that determines the strength at which the feature importance gets translated into differential feature sampling. We refer to this approach as “Pretrained GRF”. A related approach is to use $m(X_{i})$ for selecting/discarding features based on their importance. If a given feature’s importance is below a certain threshold in grf (“prescreened GRF”) it is dropped or sampled with higher probability based on its coefficients from a lasso model (“lasso-prescreened GRF”), when estimating the $\tau(X_{i})$ model. Figure 8 shows the performance of these modified CATE estimators using the DGP described in subsection 3.2 using the grf package (Tibshirani et al., 2024).
@@
Figure 8: Effect of Pretraining Strategies for Generalized Random Forests (GRF). The figure shows the performance in terms of mean squared error (MSE), normalized by the baseline GRF, in three different configurations: when using weighted feature sampling based on $m(x)$ (“Pretrained GRF”, left panel), prior variable prescreening and removal (“Prescreened GRF”, middle pane) based on importance scores, and prescreening based on the lasso (“Lasso-Prescreened GRF”, right panel). The experiment is conducted using the DGP described in subsection 3.2, with $p=150$, $n=500$ and $SNR=2$. The degree at which importance scores from $m(x)$ are translated into differential feature sampling/screening is determined by the $R$-loss.
@@
Finally, we test these models against the pretrained $R$-lasso in combination with basis expansions generated by a gradient boosted tree. To test these approaches we modified four DGPs developed by Nie and Wager (2021). We consider four setups A-D, each varying in the complexity of the treatment effect $\tau(X_{i})$ and the main effects function $\mu_{(0)}(X_{i})$. Setups A-C assume a randomized controlled trial with $e(X_{i})=0.5$, while the propensities in D are given by $e(X_{i})=1/(1+\exp(-X_{i1})+\exp(-X_{i2}))$. Covariates are generated as $X\sim\mathcal{N}(0,1)^{n\times p}$. Setup A features a non-linear baseline main effect given by $\mu_{(0)}(X_{i})=\sin(\cdot X_{i1}X_{i2})+2(X_{i3}-0.5)^{2}+X_{i4}^{3}+0.5X_{i5}^{3}$ and a linear treatment effect, defined as $\tau(X_{i})=(\sum_{j=1}^{5}X_{ij})/5$. Setup B introduces high-dimensional interactions: $\mu_{(0)}(X_{i})=\max\{0,X_{i1}^{2}+X_{i2}^{2},X_{i3}^{2}\}+\max\{0,X_{i4}^{2}+X_{i5}^{2}\}+X_{i6}^{2}+X_{i7}^{2}$ and a more complex, non-linear treatment effect: $\tau(X_{i})=X_{i1}^{2}+\log(1+\sum_{j=2}^{5}e^{X_{ij}})$. Setup C flips the structure of setup A: the baseline function is linear and the treatment effect is complex. The outcome model is $\mu_{(0)}(X_{i})=(\sum_{j=1}^{5}X_{ij})/5+0.8(X_{i6}+X_{i7})$ and the treatment effect is defined as $\tau(X_{i})=\max\{0,\sum_{j=1}^{3}X_{ij}+\max\{0,X_{i4}^{2}+X_{i5}^{2}\}+X_{i6}^{2}+X_{i7}^{2}+0.5(X_{i8}+X_{i9})$. Finally, Setup D presents a baseline outcome that combines piecewise linearity and quadratic effects: $\mu_{(0)}(X_{i})=\left[\max\{0,\sum_{j=1}^{3}X_{ij}\}+\max\{0,X_{i4}+X_{i5}\}\right]/2+X_{i6}^{2}+X_{i7}^{2}$. The treatment effect is composed of contrasting additive and subtractive components: $\tau(X_{i})=\max\{0,\sum_{j=1}^{3}X_{ij}\}-\max\{0,X_{i4}+X_{i5}\}+X_{i9}^{2}+X_{i10}^{2}$.
@@
Figure 9 shows the performance of all models including the (pretrained) $R$-lasso, (pretrained) GRF, and $R$-Boost. We further include two $R$-lasso models (with and without pretraining) that take basis expansions generated from xgboost, concatenated with the original design matrix, as input: the first model configuration uses the basis functions generated from the fit on $m(X_{i})$; the second leverages basis functions generated from fits on both $m(X_{i})$ and $\tau(X_{i})$, but shields the basis functions generated from the the latter from any pretraining penalties. We report results for all combinations of $n\in\{500,1000\}$, $p\in\{15,30\}$ and SNR $\in\{0.5,2\}$.
@@
Figure 9: Comparison of CATE Estimators in Non-linear Settings. The figure shows the performance in terms of mean squared error (MSE), normalized by the baseline $R$-lasso. The experiment is conducted using the DGPs described for setups A-D, with $p=30$, $n=1000$ and $SNR=2$. The degree of pretraining in each model is determined by the $R$-loss. The $R$-lasso models with subscripts $xgboost$ denote models that take basis functions generated from the fit on of an xgboost on $m(X_{i})$ as input. The model with superscript $(m,\tau)$ leverages basis functions generated from both fits on $m(X_{i})$ and $\tau(X_{i})$. The dashed line indicates the performance of the $R$-lasso by which all MSEs are standardized.

### v2

@@ 3.1 Simulations under Varying Shared Support Sizes
Our experiments are conducted assuming a randomized trial setting. In this first experiment, we suppose $n$ units, for each we are given a binary treatment assignment $W_{i}\sim Binom(e(X_{i}))$, with $e(X_{i})=0.5$, $p$ covariates $X_{i}\sim\mathcal{N}(0,1)^{p}$ and an outcome $Y_{i}=\mu_{i}+\varepsilon_{i}$. We define $\mu_{i}=\beta_{0}+\sum_{j=1}^{n}X_{ij}\beta_{j}+W_{i}\cdot\tau(X_{i})$, with $\tau(X_{i})=\theta_{0}+\sum_{j=1}^{n}X_{ij}\theta_{j}$, where $\beta_{j}\sim U(0.5,1)$, $\theta_{j}\sim U(1,2)$, and $\varepsilon_{i}\sim\mathcal{N}\left(0,\ \frac{Var(\mu)}{SNR}\right)$. We then introduce some sparsity and limit the size of the active support for $m(X_{i})$ and $\tau(X_{i})$ to $p/3$, respectively. We start with a fully disjoint support, which we gradually increase until the support is fully shared, i.e., equal to $p/3$. We evaluate performance in terms of expected mean-squared error (MSE) $\frac{1}{n}\sum_{i=1}^{n}[\tau(X_{i})-\hat{\tau}(X_{i})]^{2}$, which we normalize by the median MSE of the baseline $R$-lasso.
In our first experiment, we looked at the effect of pretraining the $R$-lasso in the two extreme cases, where the support is either fully disjoint or fully shared. Figure 3 shows the performance of the pretrained $R$-lasso in comparison to the baseline $R$-lasso in both settings and with different pretraining strengths (upper panel). Panel b shows the evolution of performance as the shared support increases from fully disjoint to fully joint support. In the absence of any overlap, pretraining shows no effect. With increasing shared support, the pretrained $R$-lasso achieves lower MSE, especially under strong pretraining (low $\alpha$).
@@
Figure 4: Concordance of Training Set R-Loss and Test Set MSE. The figure shows the median performance (standardized by the median $R$-lasso without pretraining) oat different pretraining strengths of the $R$-lasso in a linear setting with $n=500$, $p=150$, $SNR=2$, an active set of size $|S|=\frac{p}{3}$ and for $5000$ repetitions. The results are stratified by the size of shared support between prognostic and predictive factors: disjoint (left), partially shared (middle), and fully shared support (right panel). The minimum test set MSE and training set R-loss are concordant. In settings with stronger overlap, stronger pretraining (lower $\alpha$) is favorable. The solid line at 1.0 shows the $R$-lasso without pretraining ($\alpha=1$) for comparison. Error bars denote the interquartile range between the 25th and 75th percentiles.
@@ 3.2 Assessing Power to Detect Heterogeneity
The end goal of fitting CATE models is to discover potential subgroups or strata as defined by our predictor variables $X_{i}$ that have average treatment effects that differ meaningfully from the overall population average. Quantifying their ability to achieve this goal provides valuable information. Recently proposed metrics useful for this task take inspiration from predictive analysis tools such as the ROC curve and include Chernozhukov et al. (2018); Radcliffe (2007); Sverdrup et al. (2024); Imai and Li (2023); Yadlowsky et al. (2025). An intuitive approach to operationalize this is to take inspiration from classical machine learning, where we define a training set for fitting a CATE model, and evaluate its performance on a held-out test set. We can then rank individuals from highest to lowest predicted CATE. In the presence of heterogeneity, this should yield a decreasing quantile average treatment effect as we move from higher to lower quantiles. One way to operationalize this CATE-based ranking is to use the targeting operator characteristic (TOC) curve (Yadlowsky et al., 2025), defined as
@@
The area under the TOC (AUTOC) is a numeric measure of heterogeneity and allows us to evaluate whether and how well our CATE model detects heterogeneity against a sharp null hypothesis of no heterogeneity detected. One limitation of this approach is that train-test splits can be costly and suffer from random split sensitivity, especially in settings with small sample sizes. To overcome this limitation, we adopt a sequential cross-fold approach (Wager, 2024).
Using the TOC curve, we can assess the power at which our CATE model detects heterogeneity. We hypothesized that in the presence of shared support, the pretrained $R$-lasso should be more powerful in detecting heterogeneity than without pretraining. For this experiment, we slightly tweak our DGP to a more challenging setting and reduce the support size to $p/10$, half of which is shared between $m(X_{i})$ and $\tau(X_{i})$, and the rest is disjoint. We further introduce correlation among our covariates $X\sim\mathcal{N}_{p}(\mathbf{0},\Sigma),\text{where }\Sigma\text{ is a Toeplitz matrix with entries }\Sigma_{ij}=0.5^{|i-j|}$ and reduce the SNR to 0.5. Figure 5 shows the TOC curve in two configurations: fully disjoint support and partially shared support. In the presence of overlap, the pretrained $R$-lasso shows a superior AUTOC as well as high power in detecting heterogeneity as we vary the sample size $n$ and the number of covariates $p$ (Figure 6).
@@ 3.3 Sparsity and Reliable Support Recovery: uniR-Lasso
As demonstrated in the previous section, the effectiveness of our pretraining strategy depends both on the presence of overlap and the $m(X_{i})$ model’s ability to accurately detect it. The lasso has been shown to be less reliable for this task in the presence of highly correlated features and false positives in the active set (Su et al., 2017). To address this concern, Chatterjee et al. (2025) recently developed a univariate-guided sparse regression termed “UniLasso” that achieves higher sparsity by constraining the coefficients to have the same sign and reflecting their absolute magnitudes in univariate regression.
Sparsity is also a desirable property in our setting as it will result in more covariates being penalized. Specifically, it will result in stronger, more precise pretraining when fitting the model for $\tau(X_{i})$. In addition, this model also achieves higher interpretability in high-dimensional settings with low signal, which might be desirable in many settings. As highlighted by Candes et al. (2018), identifying a subset of relevant covariates, while controlling false discovery rates (FDR), is highly valuable in many areas (Storey and Tibshirani, 2003). For example, in biology, the reliable identification of a single driver of heterogeneity can help discover a larger regulatory network that modifies treatment effects.
@@ 4 Leveraging Side Information in Non-Linear Settings
In many applications, the mean outcome and CATE functions might involve complex interactions and non-linearities that the linear lasso outlined in section 3 may not capture. A straightforward solution to continue using the lasso without modifying Algorithm 1 is to substitute the feature matrix $X_{i}$ with $\psi(X_{i})$ where $\psi:\mathcal{X}\mapsto\mathbb{R}^{d}$ is some some featurization that maps the covariates to a richer $d$-dimensional space. $\psi$ could, for example, be a collection of polynomial basis functions. With a rich set of basis functions, this approach can sufficiently capture complicated CATE or mean outcome functions (e.g., Chen (2007)). While a wide variety of basis expansions exist and the optimal choice is often application-specific (Friedman, 1991), here we concentrate on gradient boosted trees to generate these basis expansions (Chen and Guestrin, 2016; Friedman, 2001) and use the statistical software package xgboost (Chen and Guestrin, 2016). Gradient boosting is an additive model that fits a sequence of weighted trees or weak learners. We explore its use in the $R$-learner for $M=100$ and $500$ boosting rounds and a tree depth of 1 (stump). This results in a total of $M$ trees (and thus new features), each of which performs a single split. As we increase the tree depth and number of boosting rounds, the transformed feature space is able to capture more complex relations, but also suffers from higher dimensions. To perform hyperparameter tuning, i.e., select the strength of pretraining and the number of boosting rounds, we use the $R$-loss. By concatenating this new feature matrix with our original one, the lasso can now pick from the original and transformed features. This is particularly useful in partially linear settings, where estimating $m(X_{i})$ and $\tau(X_{i})$ might require a model capable of handling both linear and non-linear components simultaneously. Depending on whether we expect $m(X_{i})$ and $\tau(X_{i})$ to differ substantially in their functional forms, we can generate distinct sets of basis functions tailored to each prediction task in the $R$-learner.
@@
The lasso solution to sharing side information leverages the lasso’s adaptive regularization feature, but similar logic can be applied to other learners, such as random forests. Causal forests (Athey et al., 2019) are forest-based instantiations of the $R$-learner that uses regression forests for $m(X_{i})$ and $e(X_{i})$, and then fits $\tau(X_{i})$ using a forest-weighted version of (3) designed to express heterogeneity in the CATE. As trees rely on split criteria rather than shrinking coefficients as in the lasso, we require a new strategy that allows us to leverage side information. An effective way to improve random forests is to use adaptive feature-weighting (e.g., Basu et al. (2018)). Adapted to our setting, this procedure consists of fitting a random forest for $m(X_{i})$, deriving the feature importance score for each covariate, and, finally, sampling each feature in $\tau(X_{i})$ with a probability that reflects its importance in $m(X_{i})$. As for the lasso, we can introduce a hyperparameter that determines the strength at which the feature importance gets translated into differential feature sampling. We refer to this approach as “Pretrained GRF”. A related approach is to use $m(X_{i})$ for selecting/discarding features based on their importance. If a given feature’s importance is below a certain threshold in grf (“prescreened GRF”), it is dropped or sampled with higher probability based on its coefficients from a lasso model (“lasso-prescreened GRF”), when estimating the $\tau(X_{i})$ model. Figure 8 shows the performance of these modified CATE estimators using the DGP described in subsection 3.2 using the grf package (Tibshirani et al., 2024).
@@
Figure 8: Effect of Pretraining Strategies for Generalized Random Forests (GRF). The figure shows the performance in terms of mean squared error (MSE), normalized by the baseline GRF, in three different configurations: when using weighted feature sampling based on $m(x)$ (“Pretrained GRF”, left panel), prior variable prescreening and removal (“Prescreened GRF”, middle panel) based on importance scores, and prescreening based on the lasso (“Lasso-Prescreened GRF”, right panel). The experiment is conducted using the DGP described in subsection 3.2, with $p=150$, $n=500$, and $SNR=2$. The degree to which importance scores from $m(x)$ are translated into differential feature sampling/screening is determined by the $R$-loss.
@@
Finally, we test these models against the pretrained $R$-lasso in combination with basis expansions generated by a gradient boosted tree. To test these approaches, we modified four DGPs developed by Nie and Wager (2021). We consider four setups A-D, each varying in the complexity of the treatment effect $\tau(X_{i})$ and the main effects function $\mu_{(0)}(X_{i})$. Setups A-C assume a randomized controlled trial with $e(X_{i})=0.5$, while the propensities in D are given by $e(X_{i})=1/(1+\exp(-X_{i1})+\exp(-X_{i2}))$. Covariates are generated as $X\sim\mathcal{N}(0,1)^{n\times p}$. Setup A features a non-linear baseline main effect given by $\mu_{(0)}(X_{i})=\sin(\cdot X_{i1}X_{i2})+2(X_{i3}-0.5)^{2}+X_{i4}^{3}+0.5X_{i5}^{3}$ and a linear treatment effect, defined as $\tau(X_{i})=(\sum_{j=1}^{5}X_{ij})/5$. Setup B introduces high-dimensional interactions: $\mu_{(0)}(X_{i})=\max\{0,X_{i1}^{2}+X_{i2}^{2},X_{i3}^{2}\}+\max\{0,X_{i4}^{2}+X_{i5}^{2}\}+X_{i6}^{2}+X_{i7}^{2}$ and a more complex, non-linear treatment effect: $\tau(X_{i})=X_{i1}^{2}+\log(1+\sum_{j=2}^{5}e^{X_{ij}})$. Setup C flips the structure of setup A: the baseline function is linear, and the treatment effect is complex. The outcome model is $\mu_{(0)}(X_{i})=(\sum_{j=1}^{5}X_{ij})/5+0.8(X_{i6}+X_{i7})$ and the treatment effect is defined as $\tau(X_{i})=\max\{0,\sum_{j=1}^{3}X_{ij}+\max\{0,X_{i4}^{2}+X_{i5}^{2}\}+X_{i6}^{2}+X_{i7}^{2}+0.5(X_{i8}+X_{i9})$. Finally, Setup D presents a baseline outcome that combines piecewise linearity and quadratic effects: $\mu_{(0)}(X_{i})=\left[\max\{0,\sum_{j=1}^{3}X_{ij}\}+\max\{0,X_{i4}+X_{i5}\}\right]/2+X_{i6}^{2}+X_{i7}^{2}$. The treatment effect is composed of contrasting additive and subtractive components: $\tau(X_{i})=\max\{0,\sum_{j=1}^{3}X_{ij}\}-\max\{0,X_{i4}+X_{i5}\}+X_{i9}^{2}+X_{i10}^{2}$.
@@
Figure 9 shows the performance of all models, including the (pretrained) $R$-lasso, (pretrained) GRF, and $R$-Boost. We further include two $R$-lasso models (with and without pretraining) that take basis expansions generated from xgboost, concatenated with the original design matrix, as input: the first model configuration uses the basis functions generated from the fit on $m(X_{i})$; the second leverages basis functions generated from fits on both $m(X_{i})$ and $\tau(X_{i})$, but shields the basis functions generated from the latter from any pretraining penalties. We report results for all combinations of $n\in\{500,1000\}$, $p\in\{15,30\}$ and SNR $\in\{0.5,2\}$.
@@
Figure 9: Comparison of CATE Estimators in Non-linear Settings. The figure shows the performance in terms of mean squared error (MSE), normalized by the baseline $R$-lasso. The experiment is conducted using the DGPs described for setups A-D, with $p=30$, $n=1000$, and $SNR=2$. The degree of pretraining in each model is determined by the $R$-loss. The $R$-lasso models with subscripts $xgboost$ denote models that take basis functions generated from the fit of an xgboost on $m(X_{i})$ as input. The model with superscript $(m,\tau)$ leverages basis functions generated from both fits on $m(X_{i})$ and $\tau(X_{i})$. The dashed line indicates the performance of the $R$-lasso by which all MSEs are standardized.